# FeTa: A ferrimagnetic quantum anomalous Hall insulator

### v2

@@ Abstract
We theoretically propose that the van der Waals layered ternary transition metal chalcogenide FeTa ( S, Se, Te) is a new family of ferrimagnetic quantum anomalous Hall insulators with sizable bulk gap and Chern number . The magnetic ordering originates predominantly from the Fe atoms, where strong ferromagnetic exchange interactions between them induce magnetic moments on the Ta sites, yielding a collinear ferrimagnetic ground state. The large topological gap arises from the deep –-type band inversion between spin-down Ta and orbitals near the Fermi level—a mechanism unique to -orbital systems. Remarkably, the Curie temperature of monolayer FeTa is predicted to significantly exceed that of monolayer MnBi2Te4. Furthermore, both the Curie temperature and topological gap scale positively with the spin-orbit coupling strength of the electrons, suggesting a common physical origin. These findings, if realized experimentally, could open new avenues for the research and application of topological quantum physics.
@@
Introduction—The discovery of the quantum anomalous Hall (QAH) effect marked a milestone in the study of topological quantum states [1, 2, 3, 4]. These states exhibit an insulating bulk characterized by a finite Chern number and support gapless chiral edge modes [5, 6, 7]. The QAH phase provides a versatile platform for realizing exotic phenomena such as Majorana edge modes [8, 9, 10, 11, 12] and the quantized topological magnetoelectric effect [13, 14, 15, 16, 17], with potential for dissipationless electronic applications [18, 19]. Experimental realizations of the QAH effect span a range of material systems [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34], yet all occur at liquid helium temperatures. This low critical temperature remains a major barrier to practical use, including in metrological standards [35, 36, 37]. Identifying new QAH insulators [38, 39, 40, 41, 42, 43, 44, 45, 46] with higher transition temperatures and larger bulk gaps is therefore a central challenge in topological materials research.
@@
The QAH effect originates from spin-polarized band inversion [47], where achieving high transition temperatures and large bulk gaps necessitates a delicate interplay between two seemingly conflicting requirements: strong spin-orbit coupling (SOC) and stable ferromagnetism. From a material design perspective, strong SOC favors heavy elements, whereas ferromagnetism typically emerges in metallic systems composed of lighter element. This incompatibility has not been fully resolved in existing QAH systems. In magnetic topological insulators (TIs), the inhomogeneities drastically suppress the exchange gap by several orders of magnitude [48, 49, 50, 51, 52, 53]; moiré systems are intrinsically limited by their low energy scales [29, 30, 31]; and rhombohedral graphene suffers from a weak proximitized Ising SOC [54]. Therefore, finding stoichiometric 2D magnetic materials for the QAH effect—ideally in monolayer form with versatile tunability—is highly desired.
@@
A promising route toward high-temperature QAH insulators lies in leveraging correlated electrons, which can simultaneously support robust magnetism and nontrivial band topology. FeSe and its derivatives are particularly compelling in this context, as electron doping via lithium intercalation has revealed the coexistence of superconductivity, itinerant ferromagnetism, and topology [59, 60, 61, 62, 63, 64, 65]. Motivated by these developments, we perform density functional theory (DFT) calculations and tight-binding modeling to demonstrate that monolayer FeTa ( = S, Se, Te) emerges as a promising candidate for high-temperature QAH insulators. All compounds exhibit a ferrimagnetic ground state with Chern number and exceptionally large bulk gaps. Notably, the magnetic ordering originates primarily from the Fe atoms, while the topology arises from the spin-polarized band inversion involving the Ta and orbitals. This mechanism, where both magnetism and topology emerge cooperatively from -orbital physics, suggests a broadly applicable design principle for engineering QAH phases in correlated materials.
@@
Structure and magnetic properties—The monolayer FeTa crystallizes in a tetragonal lattice structure with the space group - (No. 115). As shown in Fig. 1(a), each primitive cell includes three atomic layers—one FeTa and two layers—where each Fe or Ta atom is coordinated by four atoms forming a distorted edge-sharing tetrahedron. This structure can be viewed as a half-substituted derivative of Fe with space group , in which one of the two inequivalent Fe atoms in the unit cell is replaced by a Ta atom. Consequently, inversion symmetry () is broken, but the combined symmetry remains preserved [Fig. 1(b)]. The fully optimized lattice constants are listed in Table 1. The dynamical and thermal stability of monolayer FeTa are confirmed by first-principles phonon calculations and molecular dynamics simulations [73], respectively. The van der Waals nature of these materials suggests that monolayers could be experimentally exfoliated from bulk samples once synthesized.
@@
The magnetic properties of monolayer FeTa are markedly altered by Ta substitution. First-principles calculations reveal that all compounds listed in Table 1 exhibit a robust collinear ferrimagnetic ground state with an out-of-plane easy axis, in contrast to monolayer FeTe, which is observed to have a stripe antiferromagnetic (AFM) ground state [74]. The total magnetic moment per unit cell is , where each Fe atom possesses sizable local moments (2.7), and Ta atom exhibits antiparallel moments (1.4). The magnetic ordering is predominantly governed by the Fe atoms, while the moments of Ta atoms are induced via strong molecular fields generated by neighboring Fe atoms [73]. This ferrimagnet effectively behaves as a ferromagnet, with the net magnetization aligned along the Fe moment direction. As a result, the system avoids a compensation temperature, consistent with Néel’s molecular field theory [75, 76].
@@
Collectively, these exchange interactions stabilize a ferrimagnetic ground state in monolayer FeTa, featuring strong FM coupling between Fe atoms and antiparallel alignment of the induced moments on Ta sites. The system exhibits exceptionally large magnetocrystalline anisotropy energy (MAE), exceeding that of known 2D magnets such as Fe3GeTe2 [78]. As listed in Table 1, the large positive MAE indicates a robust out-of-plane easy axis. Notably, the predicted Curie temperature for monolayer FeTa significantly surpasses that of MnBi2Te4.
Electronic structures— Figures 2(d) and 3(a) show the electronic structure of monolayer FeTaSe2 with and without SOC, respectively. In the absence of SOC [Fig. 3(a)], four spin-polarized Dirac points emerge along the -X(-Y) direction, which are protected by mirror symmetry (). The -orbitals-resolved analysis reveals that these Dirac cones originate from the band crossings between the Ta and orbitals. Upon inclusion of SOC, the out-of-plane ferrimagnetic ordering breaks both and , opening a finite gap at each Dirac point. Each gapped Dirac cone contributes a quantized Berry phase of , yielding a total Berry phase of due to the fourfold valley degeneracy. This results in a QAH conductance of [Fig. 2(f)], corresponding to a Chern number . The nontrivial topology is further confirmed by the presence of two chiral edge modes traversing the bulk gap in the edge local density of states [Fig. 2(e)]. The sign of aligns with the direction of the Fe minority spins. Substituting Se with isovalent chalcogen atoms S or Te generates similar band structure and preserves the overall band topology and Chern number, as illustrated in Figures 2(a-c) and 2(g-i). Notably, the Dirac gap increases from S to Te, reflecting the enhancement of SOC strength with heavier chalcogen elements.
@@
In FeTa, both the topology and magnetism originate from the orbitals of transition metal elements. To highlight their shared origin, Fig. 3(c) presents the dependence of the Dirac gap and Curie temperature on the SOC strength in monolayer FeTaSe2. As the SOC increases, both the topological gap and exhibit a clear positive correlation, indicating that SOC not only drives the nontrivial band topology but also enhances the magnetic stability. This concurrent enhancement emphasizes a key advantage of -orbital systems: the intrinsic coupling between magnetism and topology. Such synergy offers an effective strategy for realizing high-temperature QAH insulators with sizable topological gaps.
Model and origin of topology—While the band structure near the Fermi level might suggest that the band topology arises from an –-type band inversion at the X point—supported by the opposite eigenvalues of the irreducible representations X1 and X4 [Fig. 3(d)] and their potential to yield a high Chern number () via valley multiplicity—a more detailed orbital analysis reveals otherwise. The -orbital band projections show that X1 originates from the unoccupied Fe band, which does not participate in the band inversion. Instead, the nontrivial topology stems from the band inversion between the Ta and orbitals at the point, involving an exchange of irreducible representations and [73]. Given the considerable depth of this inversion, its associated features are not readily apparent near the Fermi level.
To elucidate the origin of the band topology, we construct a minimal tight-binding model that incorporates the Ta , and Fe orbitals. This model, detailed in the Supplemental Material [73], includes the nearest- and next-nearest-neighbor hopping terms as well as SOC. As shown in Figs. 4(a) and 4(b), it accurately reproduces the DFT band structure [Fig. 3(a)] and the irreducible representations at high-symmetry points (listed in Table 2). Importantly, while the model captures the apparent –-like features around the X point via inter-orbital hybridization, it demonstrates that the topological band structure is fundamentally driven by the –-type inversion between the Ta and orbitals at point.
@@
Discussions—The realization of QAH state requires a delicate interplay between topology and magnetism. In FeTa, this balance is achieved through distinct yet cooperative mechanisms: the nontrivial band topology originates from SOC-driven inversion within Ta orbitals, while magnetism arises primarily from Fe electrons. Strong FM exchange between Fe atoms stabilizes local moments, inducing spin polarization on Ta sites and enabling polarized band inversion in the ferrimagnetic state. A crucial factor is the electron transfer from the less electronegative Ta to Fe, which establishes the desired Fe valence state. In contrast, LixFeSe is limited by chemically active and weakly bound Li+ ions, which cannot reach full intercalation () with current techniques [62], leading to disruptions in the topological electronic structure [41]. By comparison, Ta in FeTa forms strong and stable bonds with the ligands, thereby preserving the delicate electronic and magnetic environment essential for realizing the QAH state. This targeted substitution approach—simultaneously tuning topology, magnetism, and charge transfer—offers a general strategy for designing multifunctional materials that cohost topological, magnetic, and superconducting phases within a single compound.
FeTa constitutes the first example of a QAH insulator in a stoichiometric ferrimagnet. These materials simultaneously support spin transport and quantized edge charge currents, making them highly attractive for dissipationless electronics, spintronics, quantum computation, and magnetic tunnel junctions. The coexistence of topology and ferrimagnetism defines a promising new research direction. Owing to their inequivalent magnetic sublattices, ferrimagnets can potentially offer the combined advantages of both ferromagnets and antiferromagnets [79]: net magnetization is readily controllable via external fields, while spin dynamics remain fast, akin to AFM systems. It is intriguing to find compensated ferrimagnetic QAH insulators, in which the topological quantization efffect can be realized without any net magnetization and be controlled via fast magnetization reversal [80]. Moreover, the lattice constants of FeTa closely match those of Fe superconductors, facilitating the fabrication of FeTa/Fe heterostructures. The resulting superconducting proximity effect with reduced net magnetization offers a viable platform for realizing chiral topological superconductivity and Majorana fermions [10, 11, 12].
In summary, we identify a class of large-gap QAH insulators with Chern number , where magnetism and topology emerge cooperatively from orbital electrons alone. In contrast to earlier generations of magnetic TIs—such as magnetically doped TI and MnBi2Te4—the FeTa family uniquely combines structural simplicity, intrinsic magnetism, and topological nontriviality. We anticipate that these materials will serve as a next-generation platform for advancing research in topological quantum matter and spintronics.
@@ Acknowledgements.
Acknowledgments—We thank J. Gao, T. Yu, J. Xiao, and Y. Zhang for valuable discussions. This work is supported by the Natural Science Foundation of China through Grants No. 12350404 and No. 12174066, the Innovation Program for Quantum Science and Technology through Grant No. 2021ZD0302600, the Science and Technology Commission of Shanghai Municipality under Grants No. 23JC1400600, No. 24LZ1400100 and No. 2019SHZDZX01. Y.J. acknowledges additional support from the China Postdoctoral Science Foundation under Grants No. GZC20240302 and No. 2024M760488.References
@@
- Hu et al. [2024] Y. Hu, F. Meng, H. Lei, Q.-K. Xue, and D. Zhang, Possible spin-polarized cooper pairing in high temperature fese superconductor (2024), arXiv:2405.10482 [cond-mat.supr-con] .

### v3

@@ Abstract
We propose that the van der Waals layered ternary transition metal chalcogenides FeTa ( S, Se, Te) as a new family of ferrimagnetic quantum anomalous Hall insulators with a sizable bulk gap and high Chern number . The magnetic order arises primarily from Fe atoms, whose strong ferromagnetic exchange induces moments on neighboring Ta sites. The large topological gap originates from a deep –-type band inversion between spin-down Ta and orbitals near the Fermi level—a mechanism unique to -orbital systems. Remarkably, the Curie temperature of monolayer FeTa is predicted to significantly exceed that of monolayer MnBi2Te4. Furthermore, both the Curie temperature and topological gap scale positively with the spin-orbit coupling strength of the electrons, suggesting a common physical origin. Owing to its structural similarity to Fe superconductors and reduced net magnetization, FeTa further offers a promising platform for realizing chiral topological superconductivity. These findings, if realized experimentally, could open new avenues for the research and application of topological quantum physics.
@@
Introduction—The discovery of the quantum anomalous Hall (QAH) effect marked a milestone in the study of topological quantum states [1, 2, 3, 4]. These states exhibit an insulating bulk characterized by a finite Chern number and support gapless chiral edge modes [5, 6, 7]. The QAH phase provides a versatile platform for realizing exotic phenomena such as Majorana edge modes [8, 9, 10, 11, 12] and the quantized topological magnetoelectric effect [13, 14, 15, 16, 17], with potential for dissipationless electronic applications [18, 19]. Experimental realizations of the QAH effect span a range of material systems [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35], yet all occur at liquid helium temperatures. This low critical temperature remains a major barrier to practical use, including in metrological standards [36, 37, 38]. Identifying new QAH insulators [39, 40, 41, 42, 43, 44, 45, 46, 47] with higher transition temperatures and larger bulk gaps is therefore a central challenge in topological materials research.
@@
The QAH effect originates from spin-polarized band inversion [48], where achieving high transition temperatures and large gaps requires a delicate interplay between two seemingly conflicting requirements: strong spin-orbit coupling (SOC)—favored by heavy elements and stable ferromagnetism—typically emerges in metallic systems composed of lighter element. This incompatibility has not been resolved in existing QAH systems. In magnetic topological insulators (TIs), the inhomogeneities drastically suppress the exchange gap by several orders of magnitude [49, 50, 51, 52, 53, 54]; moiré systems are intrinsically limited by their low energy scales [30, 31, 32]; and rhombohedral graphene suffers from a weak proximitized Ising SOC [55]. Therefore, finding stoichiometric 2D magnetic materials for the QAH effect—ideally in monolayer form with versatile tunability—is highly desired.
@@
A promising route toward high-temperature QAH insulators lies in leveraging correlated electrons, which can simultaneously support robust magnetism and nontrivial band topology. FeSe and related compounds exemplify this potential, as electron doping via lithium intercalation has revealed the coexistence of superconductivity, itinerant ferromagnetism, and topology [56, 57, 58, 59, 60, 61, 62]. Motivated by these advances, we perform density functional theory (DFT) calculations and tight-binding modeling to show that monolayer FeTa ( = S, Se, Te) is a compelling candidate for high-temperature QAH insulators. Structural relaxations and magnetic properties are obtained within [63, 64, 65], while the electronic structure and MAE are confirmed using the Heyd-Scuseria-Ernzerhof hybrid functional [66]. All members of the FeTa family exhibit a ferrimagnetic ground state with high Chern number and exceptionally large topological gaps. The gap originates from a deep - band inversion between Ta and orbitals, while the strong ferrimagnetic order is driven by Fe atoms. This cooperative mechanism—where both magnetism and topology emerge from -orbital physics—highlights a general design principle for engineering robust QAH phases in correlated materials.
@@
Structure and magnetic properties—The monolayer FeTa crystallizes in a tetragonal lattice structure with the space group - (No. 115). As shown in Fig. 1(a), each primitive cell includes three atomic layers—one FeTa and two layers—where each Fe or Ta atom is coordinated by four atoms forming a distorted edge-sharing tetrahedron. This structure can be viewed as a half-substituted derivative of Fe with space group , in which one of the two inequivalent Fe atoms in the unit cell is replaced by a Ta atom. Consequently, the inversion symmetry () is broken, but the combined symmetry remains preserved [Fig. 1(b)]. The fully optimized lattice constants are listed in Table 1. The dynamical and thermal stability of monolayer FeTa are confirmed by phonon calculations, formation energies and molecular dynamics simulations [74]. In reality, FeCu with the same structure have been experimentally synthesized [75, 76, 77, 78]. Meanwhile, their van der Waals layered structure implies that monolayers can be readily exfoliated from bulk crystals once synthesized.
@@
The magnetic properties of monolayer FeTa are markedly altered by Ta substitution. First-principles calculations reveal that all compounds listed in Table 1 exhibit a robust collinear ferrimagnetic ground state with an out-of-plane easy axis, in contrast to monolayer FeTe, which is observed to have a stripe antiferromagnetic (AFM) ground state [79]. The total magnetic moment per unit cell is , where each Fe atom possesses sizable local moments (), and Ta atom exhibits antiparallel moments (). The magnetic ordering is predominantly governed by the Fe atoms, while the moments of Ta atoms are induced via strong molecular fields generated by neighboring Fe atoms [74]. This ferrimagnet effectively behaves as a ferromagnet, with the net magnetization aligned along the Fe moment direction. As a result, the system avoids a compensation temperature, consistent with Néel’s molecular field theory [80, 81].
@@
Collectively, these exchange interactions stabilize a ferrimagnetic ground state in monolayer FeTa, featuring strong FM coupling between Fe atoms and antiparallel alignment of the induced moments on Ta sites. The system exhibits exceptionally large MAE, exceeding that of known 2D magnets such as Fe3GeTe2 [83]. As listed in Table 1, the large positive MAE indicates a robust out-of-plane easy axis. Notably, the predicted Curie temperature for monolayer FeTa significantly surpasses that of MnBi2Te4.
Electronic structures— Fig. 2(d) show the electronic structure of monolayer FeTaSe2 with and without SOC. In the absence of SOC, four spin-polarized Dirac points emerge along the -X(-Y) direction, which are protected by mirror symmetry (). Upon inclusion of SOC, the out-of-plane ferrimagnetic ordering breaks both and , opening a finite gap at each Dirac point. Each gapped Dirac cone contributes a quantized Berry phase of , yielding a total Berry phase of due to the fourfold valley degeneracy. This results in a QAH conductance of [Fig. 2(f)], corresponding to a Chern number . The nontrivial topology is further confirmed by the presence of two chiral edge modes traversing the bulk gap in the edge local density of states [Fig. 2(e)]. The sign of aligns with the direction of the Fe minority spins. The partial filling of high-Chern-number bands could lead to new topological states with exotic elementary excitations [84]. Substituting Se with isovalent chalcogen atoms S or Te generates similar band structures and preserves the overall band topology and Chern number, as illustrated in Figures 2(a-c) and 2(g-i). Notably, the Dirac gap increases from S to Te, reflecting the enhancement of SOC strength with heavier chalcogen elements.
@@
In FeTa, both the topology and magnetism originate from the orbitals of transition metal elements. To highlight their shared origin, Fig. 3(c) presents the dependence of the Dirac gap and Curie temperature on the SOC strength in monolayer FeTaSe2. As the SOC increases, both the topological gap and exhibit a clear positive correlation, indicating that SOC not only drives the nontrivial band topology but also enhances magnetic stability. This concurrent enhancement emphasizes a key advantage of -orbital systems: the intrinsic coupling between magnetism and topology. Such synergy offers an effective strategy for realizing high-temperature QAH insulators with sizable topological gaps.
Model and origin of topology—While the band structure near the Fermi level might suggest that the band topology arises from an –-type band inversion at the X point—supported by the opposite eigenvalues of the irreducible representations X1 and X4 [Fig. 3(d)] and their potential to yield a high Chern number () via valley multiplicity—a more detailed orbital analysis reveals otherwise. The -orbital band projections and elementary band representation analysis in Fig. 3(a) show that X1 originates from the unoccupied Fe band, which does not participate in the band inversion. Instead, the nontrivial topology stems from the band inversion between the Ta and orbitals at the point, involving an exchange of irreducible representations and [74]. Given the considerable depth of this inversion, its associated features are not readily apparent near the Fermi level.
To elucidate the origin of the band topology, we construct a minimal tight-binding model that incorporates the Ta , and Fe orbitals. This model, detailed in the Supplemental Material [74], includes the nearest- and next-nearest-neighbor hopping terms as well as SOC. As shown in Figures 4(a) and 4(b), it accurately reproduces the DFT band structure [Fig. 3(a)] and the irreducible representations at high-symmetry points (listed in Table 2). Importantly, while the model captures the apparent –-like features around the X point via inter-orbital hybridization, it demonstrates that the topological band structure is fundamentally driven by the –-type band inversion between the Ta and orbitals at point.
@@
Discussions—The high-temperature QAH states in FeTa arise from cooperative mechanisms of topology and magnetism entirely within orbitals. The nontrivial band topology is driven by SOC-induced inversion of Ta orbitals, while robust magnetism originates from Fe electrons. The Fe–Ta -orbital hybridization is substantial [74]. In sharp contrast, MnBi2Te4 derives its topology from Bi/Te orbitals at the Fermi level, while magnetism comes from Mn orbitals located far from the gap. The weak – hybridization between Mn and Te yields relatively weak exchange interactions, which not only suppress the critical temperature but also limit the size of the topological gap. This cooperative interplay of magnetism and topology arising from -orbital physics highlights a broadly applicable design principle for realizing QAH phases with enhanced gaps and elevated critical temperatures.
FeTa constitutes the first example of a QAH insulator in a stoichiometric ferrimagnet. Owing to their inequivalent magnetic sublattices, ferrimagnets can potentially offer the combined advantages of both ferromagnets and antiferromagnets [85]: net magnetization is readily controllable via external fields, while spin dynamics remain fast, akin to AFM systems. The coexistence of topology and ferrimagnetism could open a promising new research direction in the interdisciplinary fields of topological materials and spintronics. It is intriguing to find compensated ferrimagnetic QAH insulators, in which the topological quantization efffect and dissipationless spintronic devices can be realized without any net magnetization and be controlled via fast magnetization reversal [86].
FeTa also bears an intrinsic connection to superconductivity. Its lattice structure closely matches that of FeSe and related superconductors, enabling the fabrication of FeTa/Fe heterostructures. The resulting superconducting proximity effect, combined with reduced net magnetization, provides a viable platform for realizing chiral topological superconductivity and Majorana fermions [10, 11, 12]. Moreover, drawing inspiration from LixFeSe—where increasing lithium concentration drives a crossover from a nonmagnetic superconductor to a superconducting ferromagnet [57, 58, 59]—one may tune the substitution ratio in Fe2-xTa (). Such tuning could yield a rich phase diagram that may evolve from a nonmagnetic superconductor to a time-reversal-breaking superconducting magnet, and ultimately to a QAH insulator.
In summary, we identify a family of large-gap, high-temperature QAH insulators with high Chern number , in which magnetism and topology emerge cooperatively from orbital electrons. In contrast to earlier generations of magnetic TIs—such as magnetically doped TI and MnBi2Te4—the FeTa compounds uniquely combine structural simplicity, intrinsic magnetism, and nontrivial topology. We anticipate that these materials will serve as a next-generation platform for exploring topological quantum matter and advancing spintronic applications.
@@ Acknowledgements.
Acknowledgments—We thank J. Gao, T. Yu, J. Xiao, and Y. Zhang for valuable discussions. This work is supported by the Natural Science Foundation of China through Grants No. 12350404 and No. 12174066, the Innovation Program for Quantum Science and Technology through Grant No. 2021ZD0302600, the Science and Technology Commission of Shanghai Municipality under Grants No. 23JC1400600, No. 24LZ1400100 and No. 2019SHZDZX01, and is sponsored by the “Shuguang Program” supported by the Shanghai Education Development Foundation and Shanghai Municipal Education Commission. Y.J. acknowledges additional support from the China Postdoctoral Science Foundation under Grants No. GZC20240302 and No. 2024M760488. Y.J. and H.W. contributed equally to this work.References
@@
- Guo et al. [2025] Z. Guo, J. Gao, Z. Huang, D. Yue, Z. Liu, M. Luo, S. Wu, X. Chen, G. Huang, Y. Deng, et al., Quantized topological states and parity anomaly in intrinsic quantum anomalous hall insulator mnbi2te4 (2025), arXiv:2507.03342 [cond-mat.mes-hall] .
@@
- Hu et al. [2025] Y. Hu, K. Liang, J. Li, Z. Li, F. Meng, H. Lei, J. Wang, H. Wen, R. Zhang, J. Cai, J. Zhang, Y. Lu, Y. Wang, Q.-K. Xue, and D. Zhang, Lithium intercalated fese as a high-temperature superconducting ferromagnet, Nature Commun. 16, 7305 (2025).
@@
- Kradinova et al. [1993] L. V. Kradinova, A. M. Polubotko, V. V. Popov, V. D. Prochukhan, Y. V. Rud, and V. E. Skoriukin, Novel zero-gap compounds, magnetics: Cufes2 and cufete2, Semicond. Sci. Technol. 8, 1616 (1993).
- Rivas et al. [1998] A. Rivas, F. Gonzalez-Jimenez, L. D’Onofrio, E. Jaimes, M. Quintero, and J. Gonzalez, Mössbauer measurements in cufete2, Hyperfine interactions 113, 493 (1998).
- Llanos and Mujica [1995] J. Llanos and C. Mujica, Synthesis and characterization of a new copper iron telluride, J. Alloys Compd. 217, 250 (1995).
- Ga et al. [2022] Y. Ga, Q. Cui, Y. Zhu, D. Yu, L. Wang, J. Liang, and H. Yang, Anisotropic dzyaloshinskii-moriya interaction protected by crystal symmetry in two-dimensional ternary compounds, npj Comput. Mater. 8, 128 (2022).
@@
- Barkeshli and Qi [2012] M. Barkeshli and X.-L. Qi, Topological nematic states and non-abelian lattice dislocations, Phys. Rev. X 2, 031013 (2012).